OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cos pi/3 cos pi/5 + sin pi/3 sin pi/5

OpenStudy (anonymous):

use the identity: cos (a-b)=cos a cos b +sin a sin b

OpenStudy (anonymous):

but how do u use that to solve it

OpenStudy (anonymous):

in this case you have a=pi/3 and b=pi/5 lets say then cos pi/3 cos pi/5 + sin pi/3 sin pi/5=cos (pi/3 - pi/5)=cos (-2pi/15)=cos(2pi/15) you could've different answers due to periodicity..but all are same in numeric value

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

my pleasure